# Università di Firenze, Italy

### v3

@@ Abstract
Keywords:
fractional differential equations fractional integrals terminal value problems Jacobi polynomials Fractional Hamiltonian Boundary Value Methods FHBVMs
MSC:
34A08 65R20
@@ 1 Introduction
Fractional differential equations have gained more and more importance in many applications: we refer, e.g., to the classical references Di2010 ; Po1999 for an introduction.
@@
Usually, one solves fractional initial value problems, that is, initial value problems for fractional differential equations (in short, FDE-IVPs) (see, e.g. DiFoFr2005 ; Ga2015 ; Ga2018 ; LYC16 ; Lu1985 ; SOG17 ):
@@
whose solution is
@@
However, under mild hypothesis on $f$, also the FDE-TVP is well-posed FoMo2011 , and its numerical solution has been considered by many authors (see, e.g., DiUl2023 ; FoMo2011 ; G2021 ; GK2021 ; LLZ2021 ; SWB2021 ). In particular, the scalar case of (1) ($m=1$) allows using a shooting procedure coupled with the bisection method Di2015 ; FoMoRe2014 or, more recently, with the secant method DiUl2023 .
@@
The procedure takes advantage of a recently introduced method for solving FDE-IVPs, able to obtain spectrally accurate approximations BBBI2024 . This latter approach has been derived as an extension of Hamiltonian Boundary Value Methods (HBVMs), special low-rank Runge-Kutta methods originally devised for Hamiltonian problems (see, e.g., BI2016 ; BI2018 ), and later extended along several directions (see, e.g., ABI2019 ; ABI2022-1 ; ABI2023 ; BBBI2024 ; BFCIV2022 ; BI2022 ), including the numerical solution of FDEs. A main feature of HBVMs is the fact that they can gain spectrally accuracy, when approximating ODE-IVPs ABI2020 ; BIMR2019 ; BMR2019 , and such a feature has been recently extended to the FDE case BBBI2024 .
With this premise, the structure of the paper is as follows: in Section 2 we sketch the shooting-Newton procedure for solving (1), along with a corresponding simplified variant; in Section 3 we recall the main facts about the (possibly spectrally accurate) numerical solution of FDE-IVPs recently proposed in BBBI2024 , with the extension for the shooting-Newton procedure; in Section 4 we report a few numerical tests, including the case of vector problems; at last, a few conclusions are given in Section 5.
@@ 2 The shooting-Newton procedure
Proof
In fact, from (2) and (4), one has:
@@
The previous results allow us stating the shooting-Newton procedure for solving (1) sketched in Table 1, where a suitable stopping criterion has to be adopted. Moreover, the starting approximation $\rho_{0}$ for the shooting-Newton iteration has to be chosen in some way, possibly exploiting any additional information; conversely, the choice $\rho_{0}=\eta$ (i.e., the final value in (1)) can be considered, as proposed in DiUl2023 .
Table 1: Algorithm 1: the shooting-Newton procedure.
@@
${\rm and}~{}~{}\quad\Phi^{(\alpha)}(t,\rho_{\ell})=f^{\prime}(y(t,\rho_{\ell}))\Phi(t,\rho_{\ell}),\quad t\in[0,T],\qquad\Phi(0,\rho_{\ell})=I$
@@
Proof
In fact, from (9) it follows that $\rho^{*}$ is a fixed-point of the corresponding iteration function,
@@
Proof
By using the notation about the Taylor theorem exposed before, one has:
@@
Proof
In fact, in such a case, the variational problem (7) simplifies to
@@ 2.1 A simplified Newton-iteration
Table 2: Algorithm 2: the simplified shooting-Newton procedure.
@@
${\rm and~{}compute}~{}~{}\quad\hat{\Phi}(T,\rho_{\ell})\approx\Phi(T,\rho_{\ell})$
@@
For this simplified shooting-Newton procedure, the following result holds true, the proof being similar to that of Theorem 2.2.
Theorem 2.5
Assume the hypotheses of Theorem 2.2 hold true and that $\hat{\Phi}(T,\xi)$ is continuously invertible in a neighborhood of $\xi=\rho^{*}$. Further, assume that the spectral radius of the matrix
@@
However, as one may expect, in this case the results of Theorems 2.3 and 2.4 does not hold, in general. As matter of fact, only a linear convergence can be granted and, for linear problems, convergence in one iteration cannot be expected, in general.
@@ 3 Implementing the algorithm
Following the approach in BBBI2024 , let us now explain the way we compute $y(T,\rho_{\ell})$ in Algorithms 1 and 2, and $\Phi(T,\rho_{\ell})$ in Algorithm 1.444On the contrary, $\hat{\Phi}(T,\rho_{\ell})$ in Algorithm 2 is strictly problem dependent and its computation cannot be stated in general: instead, a relevant specific case will be considered in Section 3.3. From (5) and (8), we have to compute:
@@
In general BBBI2024 , for coping with possible singularities in the derivative of the vector field at the origin, we shall consider the following graded mesh,
@@ 3.1 Piecewise quasi-polynomial approximation
The previous functions are then approximated via a piecewise quasi-polynomial approximation, as described in BBBI2024 , which we here briefly recall, and generalize to the approximation of the fundamental matrix solution, too. In more detail, with reference to (19) and (22), we shall look for approximations:
@@
Following steps similar to those in (BBBI2024, , Section 2), we consider the expansion of the vector field along the orthonormal polynomial basis, w.r.t. the weight function
@@
$P_{j}(x):=\sqrt{\frac{2j+\alpha}{\alpha}}\bar{P}^{(a-1,0)}(2x-1),\qquad x\in[0,1],\qquad j=0,1,\dots.$
@@
The approximations are derived by truncating the infinite series to a finite sum with $s$ terms. Consequently, for $n=1,\dots,N$, one obtains: [^6] We refer to ABI2022 for an efficient procedure for computing the fractional integrals $I_{j}^{\alpha}P_{j}(c)$, $j=0,\dots,s-1.$
@@
$\phi_{n-1}^{\alpha}(c,\rho_{\ell})~{}=$ (29)
$\rho_{\ell}+\frac{1}{\Gamma(\alpha)}\sum_{\nu=1}^{n-1}h_{\nu}^{\alpha}\sum_{j=0}^{s-1}J_{j}^{\alpha}\left(\frac{r^{n-\nu}-1}{r-1}+cr^{n-\nu}\right)\gamma_{j}(\sigma_{\nu},\rho_{\ell}),$
having set, for $x>1$,  [^7] We refer to (BBBI2024, , Section 3.3) for the efficient computation of such integrals.
$J_{j}^{\alpha}(x):=\int_{0}^{1}(x-\tau)^{\alpha-1}P_{j}(\tau)\mathrm{d}\tau,\qquad j=0,\dots,s-1.$ (30)
@@
$\phi_{n-1}^{\alpha}(c,\rho_{\ell})=\rho_{\ell}+\frac{h^{\alpha}}{\Gamma(\alpha)}\sum_{\nu=1}^{n-1}\sum_{j=0}^{s-1}J_{j}^{\alpha}(n-\nu+c)\gamma_{j}(\sigma_{\nu},\rho_{\ell}),$ (31)
@@
It can be shown (see BBBI2024 ) that $\phi_{n-1}^{\alpha}(c,\rho_{\ell})$ is nothing but the approximation of the memory term
$G_{n-1}^{\alpha}(c,\rho_{\ell})$
$=$ $\rho_{\ell}+\frac{1}{\Gamma(\alpha)}\sum_{\nu=1}^{n-1}h_{\nu}^{\alpha}\int_{0}^{1}\left(\frac{r^{n-\nu}-1}{r-1}+cr^{n-\nu}-\tau\right)^{\alpha-1}f(y_{n}(\tau h_{n},\rho_{\ell}))\mathrm{d}\tau,$
@@
The Fourier coefficients (28) can be approximated up to machine precision by using the Gauss-Jacobi formula of order $2k$ based at the zeros of $P_{k}(x)$, $c_{1},\dots,c_{k}$, with corresponding weights $b_{1},\dots,b_{k}$, by choosing $k$ large enough. As is explained in (BBBI2024, , Section 3), this allows formulating the discrete problem for computing them as: [^8] As is usual, the function $f$, here evaluated in a (block) vector of dimension $k$, denotes the (block) vector made up by $f$ evaluated in each (block) entry of the input argument.
@@
According to BBBI2024 , we give the following definition.
@@
$\Gamma_{j}(\sigma_{n},\rho_{\ell})$ $=$ $\alpha\int_{0}^{1}(1-\tau)^{\alpha-1}P_{j}(\tau)f^{\prime}(\sigma_{n}(\tau h_{n},\rho_{\ell}))\Psi_{n}(\tau h_{n},\rho_{\ell})\mathrm{d}\tau,$ (37)
$\qquad\qquad j=0,\dots,s-1,$
@@
$\Theta_{n-1}^{\alpha}(c,\rho_{\ell})~{}=$ (38)
$I+\frac{1}{\Gamma(\alpha)}\sum_{\nu=1}^{n-1}h_{\nu}^{\alpha}\sum_{j=0}^{s-1}J_{j}^{\alpha}\left(\frac{r^{n-\nu}-1}{r-1}+cr^{n-\nu}\right)\Gamma_{j}(\sigma_{\nu},\rho_{\ell}).$
@@
$\Theta_{n-1}^{\alpha}(c,\rho_{\ell})=I+\frac{h^{\alpha}}{\Gamma(\alpha)}\sum_{\nu=1}^{n-1}\sum_{j=0}^{s-1}J_{j}^{\alpha}(n-\nu+c)\Gamma_{j}(\sigma_{\nu},\rho_{\ell}).$ (39)
@@
By choosing values of $s$, and $k\geq s$, large enough, it can be seen that the approximations (35) and (42) provided by a FHBVM$(k,s)$ method can be accurate up to machine precision. In fact, from the analysis carried out in BBBI2024 , the error in approximating (13) and (14) is proved to be bounded by
@@
Actually, this amounts to using the method as a spectrally accurate method in time, as is the case for HBVMs ABI2020 ; ABI2023 ; BIMR2019 ; BMR2019 . This kind of approximations will be considered in the implementation of the algorithm listed in Table 1 (and for the simplified version of it, listed in Table 2), which we shall use for the numerical tests reported in Section 4.
@@ 3.2 Error estimation
by virtue of the perturbation result of Theorem 2.1, one derives the estimates
@@ 4 Numerical Tests
starting from ${\bm{\gamma}}^{n,0}={\bm{0}}$, for (33), and similarly for (40). [^10] More refined nonlinear iterations will be studied elsewhere. The iterations are carried out until full machine accuracy is gained, so that we expect full machine accuracy for the computed approximation (42) to $\Phi(T,\rho_{\ell})$, as well as a corresponding fully accurate discrete solution (44).
@@
  • the first 3 problems are the same scalar test problems in (DiUl2023, , Section 5);
@@ 4.1 Example 1
$y(t)=t^{8}-3\,t^{4.15}+\frac{9}{4}\,t^{0.3}.$ (49)
In this case, we use a uniform mesh with stepsize $h=1/10$. The method converges in 4 iterations producing the following approximations:
@@ 4.2 Example 2
We refer to Ga2015ml and the accompanying software  ml.m, for an efficient Matlab© implementation of the Mittag-Leffler function.
@@ 4.3 Example 3
For solving problem (52), we use a uniform mesh with stepsize $h=20/400=1/20$. The method converges in 6 iterations, producing the following approximations:
@@ 4.4 Example 4
$y^{(0.5)}$ $=$ $\left(\begin{array}[]{rr}-3&~{}0\\ -2&~{}-1\end{array}\right)y,\qquad t\in[0,2],$
$y(2)$ $=$ $\left(\begin{array}[]{c}2\,E_{0.5}\left(-3\cdot 2^{0.5}\right)\\[2.84526pt] 2\,E_{0.5}\left(-3\cdot 2^{0.5}\right)+E_{0.5}\left(-2^{0.5}\right)\end{array}\right)\simeq\left(\begin{array}[]{r}.2591172572977875\\[2.84526pt] .5953212597441289\end{array}\right),$

### v5

# A shooting-Newton procedure for solving fractional terminal value problems

@@ Abstract
Keywords: fractional differential equations, fractional integrals, terminal value problems, Jacobi polynomials, Fractional Hamiltonian Boundary Value Methods, FHBVMs
MSC: 34A08, 65R20
@@ 1 Introduction
Fractional differential equations have gained more and more importance in many applications: we refer, e.g., to the classical references [14, 28] for an introduction.
@@
Usually, one solves fractional initial value problems, that is, initial value problems for fractional differential equations (in short, FDE-IVPs) (see, e.g. [16, 21, 22, 26, 27, 30]):
@@
whose solution is, under suitable assumptions on $f$,
@@
However, under suitable hypothesis on $f$ and $T$, also the FDE-TVP is well-posed (see, e.g., [18] for the scalar case, and [29]). Consequently, its numerical solution has been considered by many authors (see, e.g., [17, 18, 23, 24, 25, 29]). In particular, the scalar case of (1) ($m=1$) allows using a shooting procedure coupled with the bisection method [15, 19] or, more recently, with the secant method [17].
@@
The procedure takes advantage of a recently introduced method for solving FDE-IVPs, able to obtain spectrally accurate approximations [6, 8]. This latter approach has been derived as an extension of Hamiltonian Boundary Value Methods (HBVMs), special low-rank Runge-Kutta methods originally devised for Hamiltonian problems (see, e.g., [9, 10]), and later extended along several directions (see, e.g., [1, 3, 5, 6, 7, 11]), including the numerical solution of FDEs. A main feature of HBVMs is the fact that they can gain spectrally accuracy, when approximating ODE-IVPs [2, 12, 13], and such a feature has been recently extended to the FDE case [6, 8].
With this premise, the structure of the paper is as follows: in Section 2 we sketch the shooting-Newton procedure for solving (1), along with a corresponding simplified variant; in Section 3 we recall the main facts about the (possibly spectrally accurate) numerical solution of FDE-IVPs recently proposed in [6], with the extension for the shooting-Newton procedure; in Section 4 we report a few numerical tests, including the case of vector problems; at last, a few conclusions are given in Section 5.
@@ 2 The shooting-Newton procedure
Proof  In fact, from (2) and (4), one has:
@@
The previous results allow us stating the shooting-Newton procedure for solving (1) sketched in Table 1, where a suitable stopping criterion has to be adopted. Moreover, the starting approximation $\rho_{0}$ for the shooting-Newton iteration has to be chosen in some way, possibly exploiting any additional information; conversely, the choice $\rho_{0}=\eta$ (i.e., the final value in (1)) can be considered, as proposed in [17].
Table 1: Algorithm 1 – the shooting-Newton procedure.
@@
$~{}\,{\rm and}\quad\Phi^{(\alpha)}(t,\rho_{\ell})=f^{\prime}(y(t,\rho_{\ell}))\Phi(t,\rho_{\ell}),\quad t\in[0,T],\qquad\Phi(0,\rho_{\ell})=I$
@@
Proof  In fact, from (9) it follows that $\rho^{*}$ is a fixed-point of the corresponding iteration function,
@@
Proof  By using the notation about the Taylor theorem exposed before, one has:
@@
Proof  In fact, in such a case, the variational problem (7) simplifies to
@@ 2.1 A simplified Newton-iteration
Table 2: Algorithm 2 – the simplified shooting-Newton procedure.
@@
$~{}\,{\rm and~{}compute}\quad\hat{\Phi}(T,\rho_{\ell})\approx\Phi(T,\rho_{\ell})$
@@
For this simplified shooting-Newton procedure, the following result holds true, the proof being similar to that of Theorem 2.
Theorem 5
Assume the hypotheses of Theorem 2 hold true and that $\hat{\Phi}(T,\xi)$ is continuously invertible in a neighborhood of $\xi=\rho^{*}$. Further, assume that the spectral radius of the matrix
@@
However, as one may expect, in this case the results of Theorems 3 and 4 does not hold, in general. As matter of fact, only a linear convergence can be granted and, for linear problems, convergence in one iteration cannot be expected, in general.
@@ 3 Implementing the algorithm
Following the approach in [6], let us now explain the way we compute $y(T,\rho_{\ell})$ in Algorithms 1 and 2, and $\Phi(T,\rho_{\ell})$ in Algorithm 1.777On the contrary, $\hat{\Phi}(T,\rho_{\ell})$ in Algorithm 2 is strictly problem dependent and its computation cannot be stated in general: however, a relevant specific case will be considered in Section 3.3. From (5) and (8), we have to compute:
@@
In general [6], for coping with possible singularities in the derivative of the vector field at the origin, we shall consider the following graded mesh,
@@ 3.1 Piecewise quasi-polynomial approximation
The previous functions are then approximated via a piecewise quasi-polynomial approximation, as described in [6], which we here briefly recall, and generalize to the approximation of the fundamental matrix solution, too. In more detail, with reference to (19) and (22), we shall look for approximations:
@@
Following steps similar to those in [6, Section 2], we consider the expansion of the vector field along the orthonormal polynomial basis, w.r.t. the weight function
@@
$P_{j}(x):=\sqrt{\frac{2j+\alpha}{\alpha}}\bar{P}_{j}^{(\alpha-1,0)}(2x-1),\qquad x\in[0,1],\qquad j=0,1,\dots.$
@@
The approximations are derived by truncating the infinite series to a finite sum with $s$ terms. Consequently, for $n=1,\dots,N$, one obtains: [^9] We refer to [4, 8] for efficient procedures for computing the fractional integrals $I_{j}^{\alpha}P_{j}(c)$, $j=0,\dots,s-1.$
@@
$\phi_{n-1}^{\alpha}(c,\rho_{\ell})=\rho_{\ell}+\sum_{\nu=1}^{n-1}h_{\nu}^{\alpha}\sum_{j=0}^{s-1}J_{j}^{\alpha}\left(\frac{r^{n-\nu}-1}{r-1}+cr^{n-\nu}\right)\gamma_{j}(\sigma_{\nu},\rho_{\ell}),$ (29)
having set, for $x>1$,  [^10] We refer to [6, 8] for the efficient computation of such integrals.
$J_{j}^{\alpha}(x):=\frac{1}{\Gamma(\alpha)}\int_{0}^{1}(x-\tau)^{\alpha-1}P_{j}(\tau)\mathrm{d}\tau,\qquad j=0,\dots,s-1.$ (30)
@@
$\phi_{n-1}^{\alpha}(c,\rho_{\ell})=\rho_{\ell}+h^{\alpha}\sum_{\nu=1}^{n-1}\sum_{j=0}^{s-1}J_{j}^{\alpha}(n-\nu+c)\gamma_{j}(\sigma_{\nu},\rho_{\ell}),$ (31)
@@
It can be shown (see [6]) that $\phi_{n-1}^{\alpha}(c,\rho_{\ell})$ is nothing but the approximation of the memory term
$G_{n-1}^{\alpha}(c,\rho_{\ell})=\rho_{\ell}+\frac{1}{\Gamma(\alpha)}\sum_{\nu=1}^{n-1}h_{\nu}^{\alpha}\int_{0}^{1}\left(\frac{r^{n-\nu}-1}{r-1}+cr^{n-\nu}-\tau\right)^{\alpha-1}f(y_{n}(\tau h_{n},\rho_{\ell}))\mathrm{d}\tau,$
@@
The Fourier coefficients (28) can be approximated up to machine precision by using the Gauss-Jacobi formula of order $2k$ based at the zeros of $P_{k}(x)$, $c_{1},\dots,c_{k}$, with corresponding weights $b_{1},\dots,b_{k}$, by choosing $k$ large enough. As is explained in [6, Section 3], this allows formulating the discrete problem for computing them as: [^11] As is usual, the function $f$, here evaluated in a (block) vector of dimension $k$, denotes the (block) vector made up by $f$ evaluated in each (block) entry of the input argument.
@@
According to [6] (see also [8]), we give the following definition.
@@
$\Gamma_{j}(\sigma_{n},\rho_{\ell})=\alpha\int_{0}^{1}(1-\tau)^{\alpha-1}P_{j}(\tau)f^{\prime}(\sigma_{n}(\tau h_{n},\rho_{\ell}))\Psi_{n}(\tau h_{n},\rho_{\ell})\mathrm{d}\tau,\qquad j=0,\dots,s-1,$ (37)
@@
$\Theta_{n-1}^{\alpha}(c,\rho_{\ell})=I+\sum_{\nu=1}^{n-1}h_{\nu}^{\alpha}\sum_{j=0}^{s-1}J_{j}^{\alpha}\left(\frac{r^{n-\nu}-1}{r-1}+cr^{n-\nu}\right)\Gamma_{j}(\sigma_{\nu},\rho_{\ell}).$ (38)
@@
$\Theta_{n-1}^{\alpha}(c,\rho_{\ell})=I+h^{\alpha}\sum_{\nu=1}^{n-1}\sum_{j=0}^{s-1}J_{j}^{\alpha}(n-\nu+c)\Gamma_{j}(\sigma_{\nu},\rho_{\ell}).$ (39)
@@
By choosing values of $s$, and $k\geq s$, large enough, it can be seen that the approximations (35) and (42) provided by a FHBVM$(k,s)$ method can be accurate up to machine precision. In fact, from the analysis carried out in [6], the error in approximating (13) and (14) is proved to be bounded by
@@
Actually, this amounts to using the method as a spectrally accurate method in time, as is the case for HBVMs [2, 5, 12, 13]. This kind of approximations will be considered in the implementation of the algorithm listed in Table 1 (and for the simplified version of it, listed in Table 2), which we shall use for the numerical tests reported in Section 4.
@@ 3.2 Error estimation
by virtue of the perturbation result of Theorem 1, one derives the estimates
@@ 4 Numerical Tests
starting from ${\bm{\gamma}}^{n,0}={\bm{0}}$, for (33), and similarly for (40). [^13] More refined nonlinear iterations are described in [8]. The iterations are carried out until full machine accuracy is gained, so that we expect full machine accuracy for the computed approximation (42) to $\Phi(T,\rho_{\ell})$, as well as a corresponding fully accurate discrete solution (44).
@@
  • the first 3 problems are the same scalar test problems in [17, Section 5];
@@ 4.1 Example 1
$y(t)=t^{8}-3\,t^{4.15}+\frac{9}{4}\,t^{0.3}.$
In this case, we use a uniform mesh with stepsize $h=1/10$. The method converges in 4 iterations producing the approximations in Table 3.
Table 3: Results for Problem (48).
@@ 4.2 Example 2
We refer to [20] and the accompanying software  ml.m, for an efficient Matlab© implementation of the Mittag-Leffler function.
@@ 4.3 Example 3
For solving problem (50), we use a uniform mesh with stepsize $h=20/400=1/20$. The method converges in 6 iterations, producing the approximations listed in Table 5.
Table 5: Results for Problem (50).
@@ 4.4 Example 4
$y^{(0.5)}$ $=$ $\left(\begin{array}[]{rr}-3&~{}0\\ -2&~{}-1\end{array}\right)y,\qquad t\in[0,2],$ (53)
$y(2)$ $=$ $\left(\begin{array}[]{c}2\,E_{0.5}\left(-3\cdot 2^{0.5}\right)\\[2.84526pt] 2\,E_{0.5}\left(-3\cdot 2^{0.5}\right)+E_{0.5}\left(-2^{0.5}\right)\end{array}\right)\simeq\left(\begin{array}[]{r}.2591172572977875\\[2.84526pt] .5953212597441289\end{array}\right),$ (58)